Circular No. 6284

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

GALACTIC CENTER W. S. Paciesas, University of Alabama, Huntsville; B. A. Harmon and

G. J. Fishman, Marshall Space Flight Center (MSFC), NASA; and S. N. Zhang and C. R. Robinson, Universities Space Research Association and MSFC, report: “BATSE occultation monitoring indicates the emergence of a new hard x-ray source, designated GRO J1744–28, in the region of the Galactic Center. The location (equinox 2000.0) is within a region whose vertices are α = 17^{h}42^{m}, δ = −27^{o}15′; 17^{h}45^{m}, −28^{o}25′; 17^{h}43^{m}, −28^{o}43′, 17^{h}47^{m}, −29^{o}46′ (90-percent confidence). Archival data show indications of the source at gradually increasing intensity for approximately the past month. However, systematic confusion with other sources in the region limits our ability to quantify the time of earliest detection. The average intensity between Jan. 1 and 7 was ∼0.58 photons cm^{−2} s^{−1} in the band 20–100 keV, with small (10 percent) daily variations. The spectrum can be approximately represented by a steep power law with photon number index −4.2. Contributions to the flux from nearby sources cannot be excluded. The new source may be the persistent counterpart to the bursting source previously reported (IAUC 6272, 6275, 6276). If so, the combination of the occultation source position with that reported in IAUC 6276 provides a more precise location for this unusual source.”

NOVA CASSIOPEIAE 1995 U. Munari, Asiago Observatory, communicates: “Spectra (range 420–

690 nm, resolution 0.04 nm) obtained with the Asiago 1.82-m telescope (+ échelle spectrograph + CCD) on Jan. 3.8 UT show a marked increase in the intensity and FWHM of the emission lines, compared with similar observations in 1995 Nov. (cf. IAUC 6259). The profiles of hydrogen Balmer and permitted Fe II lines still show a narrow and deep absorption (FWHM 65 km/s) superimposed onto a now broad emission (FWHM 650 km/s) component, shifted by ∼ −85 km/s from the latter. Forbidden lines of O III and O I are absent, and those of Fe II are quite weak and as broad as the permitted ones.”

COMET 116P/WILD 4 Total magnitude estimates: 1995 Nov. 21.17 UT, 16.3 (H. Mikuž,

Ljubljana, Slovenia, 0.36-m reflector + V filter + CCD); Dec. 19.35, 14.1 (A. Hale, Cloudcroft, NM, 0.41-m reflector; visual); 27.31, 14.0 (Hale).

1996 January 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green